i need help on finding the value a.
if the graph of the equation 2x+5y=4 passes through the point with coordinates (7,a), the value of a is?

Answer by ewatrrr(24785)

2x+5y=4
0r
y = (-2/5)x + 4/5
P(7,a)
a = (-2/5)7 + 4/5 = -14/5 + 4/5 = -10/5 = -2
a = -2